### v2

# Advancing Post-OCR Correction: A Comparative Study of Synthetic Data

Abstract
This paper explores the application of synthetic data in the post-OCR domain on multiple fronts by conducting experiments to assess the impact of data volume, augmentation, and synthetic data generation methods on model performance. Furthermore, we introduce a novel algorithm that leverages computer vision feature detection algorithms to calculate glyph similarity for constructing post-OCR synthetic data. Through experiments conducted across a variety of languages, including several low-resource ones, we demonstrate that models like ByT5 can significantly reduce Character Error Rates (CER) without the need for manually annotated data, and our proposed synthetic data generation method shows advantages over traditional methods, particularly in low-resource languages111Code and data are available at https://github.com/NikoGuan/P_OCR.
1 Introduction
Digital libraries, like the Internet Archive, offer a vast collection of historical and culturally important books in image formats, including works written in low-resource and endangered languages. However, their image-only format limits content accessibility, hindering the use of these essential resources. Therefore, Optical Character Recognition (OCR) technologies are evidently useful in this context. However, OCR outputs frequently contain errors, particularly when working with texts featuring complex styles, archaic fonts, or unconventional layouts. These errors may include character recognition mistakes, formatting issues, and hyphenation problems, which are particularly prominent when dealing with low-resource languages Ignat et al. (2022). Poor quality OCR can reduce the usefulness of these digital texts, adversely affecting downstream tasks (Linhares Pontes et al., 2019; Koudoro-Parfait et al., 2021).
Post-OCR correction is crucial for multiple reasons, including cultural heritage preservation Jarlbrink and Snickars (2017), expanding the accessibility of knowledge and information Bazzo et al. (2020), and supporting further downstream tasks in cultural analytics Stubbs (1996). Additionally, accurate historical text data is extremely important for training large language models (LLMs) Bubeck et al. (2023), which require high-quality data to improve their ability to handle complex queries, especially those involving history and culture.
Crowdsourcing has been the primary approach to acquire post-OCR training data in this domain Clematide et al. (2016); Richter et al. (2018); Maheshwari et al. (2022). While this can provide highly accurate training data, the process is often time consuming and expensive. With the advent of Transformer architecture and attention mechanisms Vaswani et al. (2017), deep learning models have emerged as the standard approach for post-OCR tasks. These models require large amounts of data for training. Thus, synthetic data has been increasingly adopted in this context. D’hondt et al. (2017); Davydkin et al. (2023); Jasonarson et al. (2023). However, most existing literature on generating synthetic data relies on additional existing data for generation, and no comprehensive comparison has been made between different methods to understand how the synthetic data generated in various ways affects post-OCR performance.
To address these issues, this paper explores the impact of data volume and data augmentation methods on the performance of post-OCR models. We examine several common methods for creating synthetic data in the post-OCR domain and propose a novel method based on feature detection algorithms from computer vision to calculate glyph similarity for synthetic data construction. We conduct experiments on eight languages, including several low-resource languages, achieving significant CER reductions ranging from 12.41% to 48.18%.
2 Related Work
Popular OCR systems include the Google Vision API OCR system Fujii et al. (2017) and the Tesseract OCR engine Smith (2007). Jatowt et al. (2019) performed statistical analysis on the types of errors commonly produced by OCR systems.
Post-OCR correction, while often overlooked, is an important NLP task. Lexical approaches to post-correction concentrate on character and word level inaccuracies, primarily employing dictionaries and heuristic rules. Bassil and Alwani (2012) exploited Google’s search suggestions for context-based corrections, circumventing the need for exhaustive dictionaries. Strategies specific to certain domains, such as those proposed by Furrer and Volk (2011), Estrella and Paliza (2014), and Kettunen (2016), highlight the necessity of tailored dictionaries for texts that possess unique features, like historical typefaces. Wemhoener et al. (2013) focused on aligning and merging outputs from various scans, including those from different editions, to rectify errors. Recent studies have framed post-OCR tasks as Seq2Seq tasks, with researchers applying both Statistical Machine Translation (SMT) and Neural Machine Translation (NMT) models (Amrhein and Clematide, 2018). Nguyen et al. (2020) and Soper et al. (2021) employed BERT Devlin et al. (2018) and BART Lewis et al. (2019) models, respectively. Maheshwari et al. (2022) conducted a comparison between pre-trained models and traditional Seq2Seq models, with their findings indicating that pre-trained models like ByT5 Xue et al. (2022) outperform the conventional models. Ramirez-Orta et al. (2022) segmented documents into character n-grams, before aggregating their corrections into the final output via majority voting Lam and Suen (1997), essentially acting as an ensemble of sequence models.
Data is fundamental to the success of deep learning, as an increasing amount of research is directed towards leveraging data-driven strategies. These strategies aim to significantly improve model performance by optimizing data usage, rather than modifying the underlying model structure Tarafdar et al. (2019); Mazumder et al. (2022). These efforts often involve generating large quantities of synthetic data Choi and Park (2023), involving data manipulation measures like filtering Koehn et al. (2020), data augmentation Shorten and Khoshgoftaar (2019); Li et al. (2022), and noise injection Izumi et al. (2003). Such synthetic data has been widely used in various NLP tasks, such as grammatical error correction Ingólfsdóttir et al. (2023), language identification Ahmadi et al. (2023), question answering Puri et al. (2020), and named entity recognition Liu et al. (2021).
For the task of text denoising, the primary method for constructing synthetic data is noise injection Izumi et al. (2003). This technique involves inserting errors into clean text to generate training data pairs. D’hondt et al. (2017) added artificial OCR errors into sentences using a random process, while Jasonarson et al. (2023) focused on the low-resource Icelandic language for post-OCR tasks. The authors extracted OCR errors from real digitised documents and inserting them into clean text in similar proportions. Grundkiewicz et al. (2019) analyzed real data to create replacement sets for each word. Davydkin et al. (2023) adopted a different approach, developing a system to generate handwritten image data which were then processed with OCR. The resulting OCR outputs and the original texts were subsequently used to train a T5 model Raffel et al. (2020) for correction purposes. Ignat et al. (2022) synthesized text image datasets by manipulating parameters, like font spacing and image saturation, then compare the original text with the OCR output text, to evaluate OCR systems’ performance on various low-resource languages, and they enhanced Machine Translation (MT) through backtranslation Sennrich et al. (2015).
Some studies have explored improving post-OCR text correction by analyzing the visual forms of characters, known as glyphs. Chen and Zhou (2023) attempting to use the CharBERT model Ma et al. (2020) for post-OCR. Their method consists of two parallel CNN encoders and a transformer decoder, taking CharBERT and glyph embedding as inputs. Amrhein and Clematide (2018)’s experiment included NMT models and glyph embedding, but it did not enhance model performance. Other research has focused on the detection of homoglyphs, with Ginsberg and Yu (2018) employing a grid method to assess the similarity of glyphs by counting the number of overlapping grids between characters. The similarity of glyphs is actually based on visual features. In the field of computer vision, feature detection and matching algorithms, such as SIFT Ng and Henikoff (2003), ORB Rublee et al. (2011), and AKAZE Alcantarilla and Solutions (2011) can extract feature points from an image and match them with those appearing in other images.
3 Methods
We now describe three common methods for generating synthetic data in the post-OCR domain, before introducing a new method based on glyph similarity in Section 3.4. Each method makes use of a clean corpus . As a common preprocessing step, we segment into multiple chunks by initially dividing sentences based on punctuation marks and then concatenating these sentences, ensuring that the total length of each chunk does not exceed a fixed limit of 230 characters, with the exception of Russian and Telugu, where the limits are 140 and 90 characters, respectively.
3.1 Method Random Injection
The random injection method D’hondt et al. (2017) generates a synthetic OCR corpus by randomly inserting errors into an existing clean corpus .
First, we filter out infrequently occurring characters in , as the corpus may inevitably contain some noise. The remaining characters are used for replacements and insertions. Next, for each chunk from , we randomly select a target error rate , which controls the level of noise to be introduced. According to the analysis by Jatowt et al. (2019), the average rate of OCR errors involving substitution, insertion, and deletion is approximately 5:1:1. Therefore, in our implementation, each character has a probability of of being replaced by a random character and a probability of of being deleted. Additionally, any two characters have a probability of of having a random character inserted between them.
3.2 Method Image Creation
Following some OCR-related studies that add noise to text images Jaderberg et al. (2014); Krishna et al. (2018); Boros et al. (2022), we also investigate the simulation of real-world correction scenarios by creating synthetic text images. Specifically, each chunk from corpus is used to create a separate image which is manipulated to add OCR-like noise. These images are then processed through a standard OCR system to obtain .
The complete procedure is as follows. Initially, for every text chunk, a random font from the set , which matches the language of corpus , is chosen. Then, the text from each chunk is converted into an image. These images are subjected to random rotations of degrees, followed by the insertion of random noise pixels. Following this, the images undergo random dilation and erosion to simulate text variability, and their resolution are randomly reduced within a range of 0 to 50%. At the end of this process, the chosen OCR system is applied for text recognition, resulting in the output .
Figure 1 shows sample synthesized images with noise generated using this process, for several different languages. Although synthetic images are used, this method simulates the OCR process under real-world conditions, the output text is derived from the OCR system, hence we consider it to be representative of authentic OCR text.The corpora and can be used to construct test sets, training sets, or to extract distributions of OCR errors.
3.3 Method Real-World Injection
Several studies have considered the generation of synthetic data through the analysis of error distributions in existing datasets. The primary approach involves embedding OCR errors into the data at rates mirroring their occurrence in real-world settings (D’hondt et al., 2017; Grundkiewicz et al., 2019; Jasonarson et al., 2023). This strategy addresses the discrepancies between existing datasets and actual application domains, offering more control over the quality of the generated text.
To obtain a realistic OCR error distribution, in this work we use additional clean corpus coming from the same domain and apply the method described in Section 3.2 to obtain OCR-processed corpus . Subsequently, the Recursive Text Alignment Scheme (RETAS) Yalniz and Manmatha (2011) is employed to perform the automatic alignment of the OCR text and the original clean text, allowing us to extract the probabilities for character substitution, deletion, and insertion. By adjusting the probability of these errors, we can control the CER of the synthetic data. For instance, if the CER between text and its corresponding OCR version is , then doubling all of the error probabilities would result in an expected CER of for the newly generated synthetic text. In practice, we set the expected CER for each chunk in to a random value to generate the synthetic data in .
3.4 Method Glyph Similarity
Given that OCR replacement errors frequently happen between characters with visually similar glyphs Jatowt et al. (2019), we can also use the information from glyph similarities to construct synthetic data. Specifically, we apply the following procedure. Firstly, we filter infrequently appearing characters from the input corpus to form a definitive character set. Based on the analysis of the ICDAR2017 Chiron et al. (2017) and ICDAR2019 Rigaud et al. (2019) datasets, 1:1 mapping errors, where a single character is incorrectly identified as another character, accounted for 87.9% of all 1: mapping errors. Here 1: mapping error refers to cases where a single character be incorrectly recognized as characters. To enhance the computational efficiency of the implementation proposed in this paper, we focus solely on 1:1 errors. Note that simulations of 1: errors could be achieved through random insertion.
For each language, we select a set of appropriate fonts , which includes a variety of historical and modern fonts, and employ a set of vision feature detection algorithms to extract and match image features. In our experiments we consider ORB Rublee et al. (2011), AKAZE Alcantarilla and Solutions (2011) and SIFT Ng and Henikoff (2003). For each character in the character set, we calculate its glyph similarity with every other character by creating their images using fonts from the font set . We then calculate their similarity under detector as:
where is the average distance between matching point pairs for character and , using font and feature detector . The degree of overlap for feature matching points is calculated using the Jaccard Index (Jaccard, 1912). This is defined as the ratio of the number of matching points relative to the combined number of feature points in and minus the number of matching points. A number of matching examples are shown in Figure 2. We see that, as glyphs become more similar, increases and decreases.
Next, we perform min-max normalization for different mapping types and characters to obtain the normalized score
where is the set of similarity scores between character and other characters using feature detector . Figure 3 presents an example of for English-language characters.
The technique of embedding OCR errors by exploiting glyph similarity shares some common aspects with the random injection method described previously. Firstly, we randomly select a target error rate for each chunk from , where each character in has a probability of to be replaced with another character. The choice of replacement is based on , with higher-weighted pairs being more likely to be chosen. After replacement, each character has a probability of to be deleted. Any two characters have a probability of of having a random character inserted between them.
4 Datasets
In this paper, we undertake post-OCR experiments involving English, Frisian, German, Icelandic, Irish, Russian, Spanish, and Telugu texts. Clean corpora for these languages were obtained as follows: data for Icelandic, Irish, and Frisian were sourced from the CC-100 corpus Conneau et al. (2019). Telugu222https://www.kaggle.com/datasets/sudalairajkumar/telugu-nlp and Russian333https://www.kaggle.com/datasets/d0rj3228/russian-literature/data datasets were obtained from Kaggle, while English, Spanish, and German were sourced from Project Gutenberg. Summary statistics for the corpora are given in Table 1. The amount of data selected simulates the real-world rarity of these languages to some extent.
Language | Length | Source |
English | 27,883,394 | Gutenberg |
Frisian | 3,426,499 | CC100 |
German | 3,758,352 | Gutenberg |
Icelandic | 3,147,864 | CC100 |
Irish | 5,090,436 | CC100 |
Russian | 6,613,093 | Kaggle |
Spanish | 7,813,245 | Gutenberg |
Telugu | 5,777,551 | Kaggle |
All training, validation, and test datasets are generated from these clean corpora. The training data is created using the four different methods described in Section 3, with the aim of comparing the synthetic data generation techniques. To implement the method presented in Section 3.3, 20% of the text from each language corpus is used as for extracting OCR error distributions. The remaining 80% of the corpus is divided in a 8:1:1 ratio to generate training, validation, and test sets.
The test datasets are generated using a method similar to that described in Section 3.2, employing the Tesseract OCR system Smith (2007) to convert synthetic images into OCR text. Since the methods described in Sections 3.2 and 3.3 involve using an OCR system to either generate OCR text directly or to extract OCR errors, the Google Vision API OCR system Fujii et al. (2017) is used when creating training and validation datasets.
5 Models
In our experiments, we compare the performance of various models which have been previously adopted for post-OCR tasks, including models that operate at the subword-level, character-level, and byte-level tokenizers. Additionally, we evaluate current SOTA models which is based on n-gram and majority voting. This comparison encompasses both pre-trained models and those trained from scratch. Training parameters are in Appendix A.
5.1 mT5
The mT5 model Xue et al. (2020) is an extension of the T5 model Raffel et al. (2020), which is pre-trained on a multilingual dataset. This model leverages the original T5’s text-to-text framework, where every natural language processing task is reframed as a text generation problem, allowing for consistent and flexible handling of a wide range of tasks across languages. The version we use in our experiments is mT5-base.
5.2 mBART
The mBART model Tang et al. (2020) is a multilingual extension of the BART architecture Lewis et al. (2019). Developed by Facebook AI, its pre-training approach involves corrupting text with an arbitrary noising function and then learning to reconstruct the original text. mBART is also pre-trained on a large corpus of text in multiple languages, making it adept at both high-resource and low-resource language translation, as well as a variety of other language processing tasks. The version we use in our experiments is mBART-large-50.
5.3 ByT5
The ByT5 model Xue et al. (2022) is designed to address the limitations of traditional subword tokenization methods by working at the byte level. This approach allows ByT5 to handle any language or writing system with Unicode representation, making it naturally multilingual and adaptable for handling a wide variety of text data. By operating on bytes, ByT5 avoids the complexities and biases associated with subword vocabularies, enabling more equitable and accurate processing across languages. ByT5 was pre-trained on the same dataset as mT5. The model demonstrates robust performance across a wide range of tasks Stankevičius et al. (2022); Jentoft (2023). The version we use in our experiments is ByT5-base.
5.4 CharBERT + Glyph Embedding
To verify the effectiveness of glyph embedding as proposed by Chen and Zhou (2023), we also conducted some experiments using their model. CharBERT Ma et al. (2020) is a char-level model. In the pre-training phase of CharBERT, one of the tasks involves text denoising which is closely aligns with the objectives of post-OCR correction tasks. Building on the original model, Chen and Zhou (2023) introduced glyph embedding to CharBERT by using a ResNet50 model He et al. (2016). This supports the extraction of visual information from characters and serves as one of the inputs to the post-OCR correction model. The structure of this model is shown in Figure 4. In our experiments we use the same structure and fine-tuning settings corresponding to the optimal configurations reported in the original work. Since the CharBERT model was pre-trained on the English Wikipedia dataset Foundation (2024), we conduct experiments only in English with this model.
5.5 Non-pre-trained Models
The method proposed by Ramirez-Orta et al. (2022) achieved SOTA results in 5 out of 9 languages on the ICDAR2019 dataset. In this study, we also compare this model with other models using the configuration: window type as n-grams, window size of 60, decoding method as beam, weighting as uniform. We refer to as the ENSEMBLE model.
Additionally, we have defined a more conventional Seq2Seq model, referred to as SCRATCH. The SCRATCH model is designed with an embedding size of 512, feed-forward network embeddings of 2048, 4 attention heads, 5 encoder layers, 5 decoder layers, a SentencePiece tokenizer Kudo and Richardson (2018), and a vocabulary size of 3000.
6 Experiments
The primary objective of our empirical analysis in this paper is to explore the key factors affecting the performance of models in post-OCR tasks, aiming to identify the most appropriate settings for such tasks. We conduct three different experiments:
Experiment 1: We investigate how data volume impacts on model performance, aiming to identify a generally applicable data augmentation setting that balances performance and training time.
Experiment 2: We seek identify the best method for constructing synthetic data and the best-performing model. Using the findings from Experiment 1, we apply augmentation and generate different datasets using the four methods from Section 3, in combination with multiple post-OCR models.
Experiment 3: We extend the best settings identified in the first two experiments to a wider set of languages for post-OCR tasks, verifying the applicability and effectiveness of these settings in a broader linguistic context.
6.1 Experiment 1
English | Icelandic | Russian | Telugu | ||||||||||||||
CER | 4.96 | 10.09 | 4.13 | 34.12 | |||||||||||||
Model | Data | 1 | 1 | 1 | 1 | ||||||||||||
mT5 | 1 | 4.56 | 4.29 | 3.77 | 3.42 | 9.68 | 9.33 | 9.02 | 8.87 | 3.78 | 3.33 | 2.99 | 2.74 | 31.44 | 29.80 | 28.35 | 27.70 |
2 | 4.32 | 4.10 | 3.52 | 3.24 | 9.44 | 9.12 | 8.83 | 8.49 | 3.63 | 3.15 | 2.63 | 2.51 | 29.83 | 28.69 | 27.83 | 26.89 | |
4 | 4.24 | 3.71 | 3.13 | 3.00 | 9.35 | 9.00 | 8.60 | 8.31 | 3.54 | 3.05 | 2.40 | 2.33 | 29.31 | 28.45 | 27.12 | 26.31 | |
8 | 4.24 | 3.68 | 3.05 | 2.93 | 9.32 | 8.90 | 8.54 | 8.29 | 3.55 | 3.02 | 2.39 | 2.36 | 29.22 | 28.44 | 27.23 | 26.32 | |
ENSEMBLE | 1 | 4.84 | 4.69 | 4.51 | 4.46 | 10.18 | 10.09 | 10.00 | 9.88 | 4.09 | 4.00 | 3.90 | 3.83 | 40.23 | 38.47 | 38.02 | 37.71 |
2 | 4.80 | 4.63 | 4.44 | 4.30 | 10.09 | 10.03 | 9.88 | 9.74 | 4.06 | 3.85 | 3.80 | 3.71 | 38.99 | 38.65 | 38.21 | 36.60 | |
4 | 4.75 | 4.53 | 4.23 | 4.19 | 10.14 | 9.94 | 9.83 | 9.64 | 4.02 | 3.72 | 3.69 | 3.52 | 38.78 | 38.17 | 37.43 | 35.66 | |
8 | 4.79 | 4.48 | 4.24 | 4.22 | 10.11 | 9.97 | 9.80 | 9.69 | 4.00 | 3.74 | 3.63 | 3.60 | 38.65 | 37.75 | 36.54 | 35.59 |
The process of generating OCR synthetic data is highly stochastic, so the same clean text can yield different synthetic text versions when OCR errors are inserted. This variability can potentially enrich the model’s learning process. In this experiment, we investigate the impact of data augmentation techniques (by replicating the clean text data multiple times to increase the volume of data) and the quantity of original clean text data used on model performance. We examine whether increasing the diversity of data via this method can enhance the model’s ability to correct OCR errors.
The experiment is conducted on English, Icelandic, Russian and Telugu texts. These languages were chosen as they include both rich-resource and low-resource languages and belong to different language families, and have different grammatical structures and character sets. This diversity makes them representative for assessing different models and methods for generating synthetic data.
To explore the impact of the original data volume, we experiment with datasets of different sizes: the full dataset size, size, size, and size, as in post-OCR correction tasks, the training data often need to be from the same domain as the text need to be fixed, so the settings of 1, , , and were designed to simulate the scenarios within different domains under one language environment with varying amounts of clean texts. And to examine the effect of data augmentation techniques, these will be further expanded to 1, 2, 4 and 8 their original sizes. This approach will yield multiple versions of dataset , and since Method is the most commonly used, we will employ it to generate the training data for this experiment. The experiments are conducted using both the pre-trained mT5-base model and ENSEMBLE model.
The results in Table 2 show that mT5 outperforms the ENSEMBLE model across the four languages. Data augmentation enhances model performance with diminishing returns – augmenting from 1 to 4 effectively lowers CER, but gains from 4 to 8 are marginal. In our experiment, training parameters for mT5 included a learning rate of 5e-4, warm-up steps of 250, batch size of 4, dropout rate of 0.2, and 6 epochs on fp32. Training times for 4 augmented datasets on an RTX 4090 were approximately 35, 5, 10 and 8 hours for English, Icelandic, Russian and Telugu, respectively. Overall, the results suggest that 4 augmentation provides sufficient improvement.
6.2 Experiment 2
Next, we explore how different methods of creating synthetic data impact on the performance of different language models. We conduct tests using several popular models: mT5-base Xue et al. (2020), ByT5-base Xue et al. (2022), mBART-large Chipman et al. (2022), SCRATCH, and ENSEMBLE. We also evaluate the approach of Chen and Zhou (2023), which combines CharBERT Ma et al. (2020) with glyph embedding, and compare these results to CharBERT without glyph embedding.
English | Icelandic | Russian | Telugu | |||||||||||||
CER | 4.96 | 10.09 | 4.13 | 34.12 | ||||||||||||
Method | ① | ② | ③ | ④ | ① | ② | ③ | ④ | ① | ② | ③ | ④ | ① | ② | ③ | ④ |
mT5 | 3.98 | 3.42 | 3.00 | 3.03 | 9.48 | 9.53 | 8.31 | 8.38 | 3.36 | 2.98 | 2.33 | 2.27 | 30.35 | 28.24 | 26.31 | 25.85 |
ByT5 | 4.02 | 3.36 | 2.96 | 3.00 | 9.42 | 9.45 | 8.39 | 8.28 | 3.34 | 3.13 | 2.34 | 2.14 | 30.21 | 28.14 | 25.98 | 25.28 |
mBART | 4.19 | 4.23 | 3.63 | 3.54 | 9.93 | 10.26 | 9.41 | 9.30 | 3.41 | 3.33 | 2.82 | 2.78 | 31.06 | 28.49 | 25.66 | 26.00 |
CharBERT | 4.12 | 3.60 | 3.39 | 3.57 | - | - | - | - | - | - | - | - | - | - | - | - |
+Glyph | 4.11 | 3.61 | 3.42 | 3.54 | - | - | - | - | - | - | - | - | - | - | - | - |
ENSEMBLE | 4.56 | 4.44 | 4.19 | 4.00 | 10.32 | 10.65 | 9.64 | 9.54 | 3.77 | 3.83 | 3.52 | 3.34 | 37.12 | 34.23 | 35.66 | 37.50 |
SCRATCH | 4.79 | 4.74 | 4.52 | 4.62 | 11.92 | 9.82 | 10.43 | 10.00 | 3.91 | 4.06 | 3.62 | 3.64 | 33.56 | 35.28 | 34.52 | 34.92 |
English | German | Irish | Icelandic | Frisian | Russian | Spanish | Telugu | |||||||||
CER | WER | CER | WER | CER | WER | CER | WER | CER | WER | CER | WER | CER | WER | CER | WER | |
OCR | 4.96 | 15.43 | 5.79 | 19.29 | 12.57 | 35.99 | 10.09 | 30.06 | 5.15 | 16.20 | 4.13 | 8.16 | 6.00 | 17.38 | 34.12 | 90.41 |
Post-OCR | 3.00 | 7.74 | 4.27 | 11.07 | 11.01 | 29.97 | 8.28 | 24.57 | 3.55 | 11.03 | 2.14 | 5.88 | 3.76 | 8.92 | 25.28 | 66.64 |
The results of Experiment 2 are shown in Table 3. It is observed that methods and generally outperform and . Of the latter two, method , which does not rely on additional data, yields slightly higher accuracy. While method allows the model to learn from real OCR outputs, it fails to correct sentences with varying degrees of CER. In our experiments, the Google Vision API OCR system was used to generate training data, and Tesseract was used for test data. Since the Vision API OCR system generally has a higher accuracy than Tesseract, models only learn to correct sentences with low CER, which is reflected in the limitations of the method .
Method proves more effective, extracting real OCR error distributions and generating training data with various CER values, but requires additional data. With a sufficient volume of data (as in the English language experiments), models trained with this method performed best. Different OCR systems, due to their unique designs, can make various kinds of errors. With insufficient additional data, Method can fail to extract a comprehensive set of OCR errors, leading to distributional shift in the data, which ultimately affects the performance of models that rely on it. In Experiment 2, Method achieved impressive results without relying on additional data. By generating its own training data with varying degrees of CER encoded through glyphs alone, it successfully trained models that perform well across diverse languages with different character sets.
Overall, we observe that pre-trained models significantly outperformed smaller Seq2Seq models trained from scratch. mT5 and ByT5 showed similar performance, with ByT5 reducing CER values by 39.5%, 17.9%, 48.2% and 25.9% on English, Icelandic, Russian and Telugu datasets, respectively. CharBERT did not benefit from glyph embedding, consistent with the findings of Amrhein and Clematide (2018).
6.3 Experiment 3
In our final experiment, we replicate the original clean text four times for augmentation. Based on our previous experiments, we select Method to generate synthetic data and train the ByT5 model for post-OCR tasks on data from a wider set of languages, so that we can evaluate the chosen configuration in a broader linguistic context.
From the results in Table 4, we see that the model performs well on the English, Russian, and Spanish tasks, with CER reductions of 39.52%, 48.18%, and 37.33%, respectively. This approach also achieves a CER reduction of 31.07% on the low-resource language Frisian. The performance is poorest in the cases of Irish and Icelandic, with only 12.41% and 17.94% reductions in CER, respectively. We attribute this to two main reasons. Firstly, the data for Irish and Icelandic (sourced from the CC100 dataset of web-crawled data) contain numerous proper nouns. These are challenging for conventional NMT models to correct due to their inability to handle out-of-vocabulary words. Secondly, although the ByT5 model was trained on a multilingual dataset, the volume of data for low-resource languages is not substantial. This disparity in data availability can affect the models’ performance on post-OCR tasks. Table 4 also indicates that the Telugu test dataset has very high CER and WER, likely due to the complexity of Telugu glyphs (Negi et al., 2001). When Telugu text becomes blurred, it can be difficult for OCR systems to recognize. Thus, post-OCR processing for Telugu results in a 25.91% reduction in CER.
Figure 5 provides a more detailed sentence-based view of CER across languages. Here we enumerate the following categories for CER changes after post-OCR processing: "Increased" indicates sentences where the CER has increased, "Decreased" indicates the CER has decreased, "Equal" means that CER remained unchanged, and "Zero" means the CER is zero after post-OCR processing (i.e., perfect matching with the ground truth). We observe that only a small number of sentences experienced an increase in CER after post-OCR processing. Specifically, 25.37% of sentences in Irish showed an increase in CER post-OCR, and only 8.66% of sentences in Russian. Notably, 59.52% of sentences in Russian had a CER of zero after post-OCR processing. For Telugu, which had the highest CER, we saw a reduction in CER for 76.18% of sentences after post-OCR processing.
7 Conclusion
Our experiments demonstrated that augmenting data by replicating clean text and constructing synthetic data based on glyph similarity significantly improves model performance for post-OCR correction. Furthermore, we proposed a novel approach to constructing post-OCR synthetic data based on glyph similarity, which outperforms traditional noise injection methods, especially in scenarios with limited data volume, as it does not require additional external data. In our experiments, pre-trained models demonstrated superior performance compared to those without pre-training, with the byte-level ByT5 model achieving the best performance. For rich-resource languages, this approach can reduce the CER by approximately 45%, while for low-resource languages the reduction varies between 12.41% and 31.07%.
Limitations
For the preferred Method ④, the time complexity of calculating the character similarity values is . As a result, the calculation of character similarities in languages with a large array of characters, such as Chinese and Japanese, can be quite time consuming.
Acknowledgements
This publication is part of a project that has received funding from (i) the European Research Council (ERC) under the Horizon 2020 research and innovation programme (Grant agreement No. 884951); (ii) Science Foundation Ireland (SFI) to the Insight Centre for Data Analytics under grant No 12/RC/2289 P2.
Appendix A Training parameters
Due to the numerous sub-experiments in this study, models were trained on both RTX 4090 24GB and A100 80GB GPUs, without selecting different training parameters for different languages.
The mT5-base, ByT5-base, mBART-large, CharBERT, and ResNet50 models were all sourced from Hugging Face. Parameters for mT5-base, ByT5-base, and mBART-large were identical: dropout rate of 0.2, learning rate of 5e-4, batch size of 4, and 6 epochs. For CharBERT-related models, the dropout rate is 0.2, learning rate is 3e-5, batch size is 8, with 12 epochs. The ENSEMBLE and SCRATCH models have the same parameters: dropout rate of 0.1, learning rate of 1e-4, batch size of 32, and 30 epochs. All models use Adam optimizer and are trained in fp32 precision. The best models are selected based on dev loss.